# A Survey on Adversarial Contention Resolution

### v3

@@ Abstract.
survey, contention resolution, backoff, wakeup problem, selection problem, multiple access channel, broadcast, algorithms, stability, conflict resolution, Ethernet, makespan, queuing theory, adversarial queuing theory
@@ 1. Introduction
Over the past decade and a half, attention has turned to the case where the arrival times of packets are far-less constrained. One common approach has been to view this problem through the lens of adversarial queuing theory (borodin2001adversarial, ), where results are parameterized by traffic characteristics such as the arrival rate and the degree of burstiness. Another major branch in the literature addresses worst-case arrival times. Arguably, deterministic approaches have been influenced by the pioneering result by Komlos and Greenberg (komlos:asymptotically, ), leading to new solutions for worst-case arrival times. A prime example is the work by De Marco and Kowalski (doi:10.1137/140982763, ), and there has been a long line of results in this area (e.g., (DBLP:conf/icdcs/MarcoKS19, ; Chlebus:2016:SWM:2882263.2882514, ; DEMARCO20171, ; de2013contention, )). In parallel, many different randomized approaches have been explored. A recent influential paper by Chang et al. (ChangJP19, ) shows how the classic ingredients of backing off/on can yield an elegant solution to worst-case arrivals. The literature in these areas is explored in Sections 3.1, 3.2 and 3.3.
@@
Scope and Goals. In this survey, we focus primarily on contention resolution in the presence of an adversary that (1) controls when participants can begin executing their contention resolution protocol, and (2) may deny access to the shared resource for periods of time; we refer to this as adversarial contention resolution. This stands in contrast to the setting where arrival times are governed by a stochastic process, which is an area of research that has been previously surveyed by Chlebus (chlebus2001randomized, ).
What motivates our focus on adversarial contention resolution? From a theoretical standpoint, it is interesting to explore what is possible in such a challenging setting, but there are also practical considerations. Network traffic does not always align with stochastic models; specifically, bursty traffic fails to be captured by well-behaved (e.g., Poisson- or Bernoulli-distributed) arrival times. Additionally, the mobility of many modern computing devices, along with connectivity challenges in wireless systems, can lead to situations where the contention for a shared resource changes rapidly over time. Such challenges are addressed by algorithms that guarantee performance results under the problem of adversarial contention resolution.
Given this motivation, our overall goal in this survey paper is to summarize the many models and performance metrics present in the literature on adversarial contention resolution, along with discussing their corresponding theoretical results.
@@ 1.1. Definitions, Terminology, Models, Metrics, and Algorithm Design
In the literature, the resource is frequently treated as a shared communication channel, and this is the convention we use throughout this survey. Such a channel is typically referred to as a multiple access channel, which is a popular model under which several fundamental problems have been examined, such as mutual exclusion (bienkowski2010dynamic, ; czyzowicz2009consensus, ), leader election (chang:exponential-jacm, ; chang:exponential, ; nakano2002, ), consensus (czyzowicz2009consensus, ), routing (chlebus2019energy, ), broadcast problem (Marco08, ; Marco10, ; ChlebusGGPR00, ; alon1991lower, ; ClementiMS03, ; DBLP:conf/podc/Kowalski05, ; chrobak2002fast, ; piotr:explicit, ; chlebus2005almost, ; kowalski2005time, ), and well-known random access protocols (kurose:computer, ). Here, for any given slot, accessing the resource corresponds to a participant listening (equivalently, receiving) or sending (equivalently, transmitting) a packet on the multiple access channel. Unless stated otherwise, a packet can be transmitted within a single slot. When a packet is sent in a slot, we say this packet succeeds if no other packet is sent in the same slot; otherwise, the packet fails.
For ease of exposition, we will often slightly abuse terminology by speaking of packets themselves taking action (e.g., sending themselves on the channel, listening on the channel). This “packet-centric” presentation allows us to avoid referring to the corresponding “participants”/“devices”/“nodes” that actually performs these actions.
@@
We discuss model assumptions that play an important role in establishing asymptotic upper and lower bounds with respect to the metrics discussed above.
@@
In contrast, in the dynamic case, packets arrive over time; this arrival model is also referred to in the literature as asynchronous (see  (DBLP:conf/podc/Kowalski05, )). The literature adopts the terminology of arrival time or activation time to indicate when the corresponding participant begins executing a contention resolution protocol in order to send this packet. These times may be governed by a stochastic process (see the survey by Chlebus (chlebus2001randomized, )), or they may be scheduled by an adversary; in this survey, we focus on the latter setting.
Channel Characteristics. When a participant sends a packet, the multiple access channel provides feedback. The most basic feedback—assumed in all the literature we survey—occurs when, upon sending its packet, a participant learns whether or not its transmission was successful.
Another type of feedback concerns whether participants can detect collisions. Almost all of the literature on contention resolution assumes one of the following two types of feedback regarding collisions. The first is collision detection (CD): in any slot, each listening participant can differentiate between the case where the number of participants transmitting is 00 (i.e., an empty slot) or greater than $1$ (i.e., a collision). The second is no collision detection (no-CD): in any slot, a listening device cannot differentiate between the slot being empty versus containing a collision.
@@
Finally, some papers consider the case where more than one channel is available to participants; this is typically referred to as a multi-channel setting. Here, typically participants may select any one (and only one) channel to send on or listen to in any give slot. In many cases, a result involving a single channel may extend to the multi-channel setting without altering the core finding; however, there are cases where the use of multiple channels can admit significantly stronger results. Typically, this occurs when multiple channels are leveraged to achieve better latency or robustness to malicious interference, and throughout this survey, we highlight such results. In practice, multiple channels are available on several communication standards, such as IEEE 802.11 (WiFi), Bluetooth, and cellular standards (e.g., 4G and 5G LTE). Therefore, there is real-world motivation to push for results that make good use of more than one channel. However, in comparison to the single-channel setting, there are fewer results in the contention-resolution literature involving multiple channels, and so our survey is weighted more towards single-channel results.
@@
For the contention resolution problem, the type of topology may impact channel feedback. With CD in a single-hop network, all participants can detect the same channel feedback in any fixed slot. However, in a multi-hop network, it is common for different sets of participants to witness different channel feedback and, therefore, potentially update their respective states differently. Given this challenge, it is not surprising that the majority of work on contention resolution has focused the single-hop setting.
@@
  • System-Size Knowledge. The exact number of the participants may be known a priori. Alternatively, an upper bound may be common knowledge. For example, in some problem formulations (see Section 3.2), only a subset of $k\leq n$ participants have a packet to send, where $n$ acts as an upper bound that is known or tightly-estimated a priori.
  • Unique Identifiers. Given $n$ participants, each may have a unique identifier (ID) taken from a set of $\Theta(n)$ possible IDs. This ID may be used by an algorithm to determine when the corresponding station attempts to send.
@@
In this section, we adopt the packet-centric terminology. In the static case, a natural performance measure is the number of slots until a set of packets succeeds; this number of slots is referred to as the makespan. We are typically interested in the set of all packets, but any deviations from this will be made clear from the context. For the static case, the ratio of the number of successful packets to the makespan is known as throughput, and this is another common notion of performance; note that throughput is a value between 00 and $1$.
For the dynamic case, we define a slot to be active, if at least one packet is present in the system during that slot. The above metrics remain valid when measured over active slots. However, at intermediate points in time, throughput may offer an overly-pessimistic measure of progress. For example, if a batch of $n$ packets arrive at the same time and start executing BEB, the throughput will be 00 for $\Omega(n/\log n)$ slots, despite the fact that the algorithm is progressing towards having many successful packet transmissions, which occurs after $\Theta(n)$ slots are executed.
To provide a more meaningful metric, Bender et al. (bender2020contention, ) propose the notion of “implicit throughput”. The implicit throughput at time $t$ is defined as $N_{t}/A_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$, and $A_{t}$ is the total number of active slots at or before time $t$. Returning to the example with a batch of $n$ packets, we note that the implicit throughput is large—specifically, $n$—in the first slot, and is $\Omega(\log n)$ over the first $\Omega(n/\log n)$ slots. The implicit throughput will continue to decrease, reaching $\Theta(1)$. Then, ultimately, the implicit throughput will be $\Theta(1/\log n)$ once all packets have succeeded, which matches the standard measure of throughput for BEB, as mentioned in Section 1. An algorithm that guarantees an implicit throughout of $\Omega(1)$ over all slots performs well, since the number of active slots so far is never asymptotically larger than the number of packets that have arrived. Furthermore, as exemplified above, at points in time where the standard notion of throughput is meaningful—such as when there are no packets in the system—the two metrics are equal.
Continuing with the dynamic case, another metric is the number of slots it takes for a participant to successfully transmit its packet starting from its activation time; this is latency. We are interested in the maximum latency over all participants. Here, we can extend throughput to be the number of packets over the maximum latency.
@@
Given the popularity of small, battery-powered wireless devices in modern networks, many contemporary results focus on energy efficiency. The energy consumed by a participant is typically measured as the number of slots that the participant accesses the channel, as sending and listening often dominate the operational costs of such devices (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). We refer to this as the number of channel accesses.
Robustness to malicious interference—often referred to as jamming—is becoming increasingly important, with several well-publicized attacks in recent years (nicholl, ; FCC:humphreys, ; marriott, ; jamming-church, ). Contention-resolution algorithms that rely on channel feedback are especially vulnerable to adversarial noise. We refer to this aspect as attack resistance and, for each relevant result, we describe the type of attack that can be tolerated.
@@
There is some common terminology used in the literature to characterize different classes of contention resolution algorithms. An obvious division is with respect to deterministic versus randomized algorithms, where the former is characterized by the absence of random bits being used in the execution. Deterministic solutions are appealing from a practical standpoint in that their guarantees hold even when the number of participants and/or packets is small, admitting no error probability. In trade, they tend to provide sub-optimal asymptotic results in many settings, compared to randomized solutions, as we shall see.
@@ 1.2. Survey Organization
We use the attributes discussed above and illustrated in Figure 1 to structure our review of the literature. In Section 2, we begin with contention-resolution algorithms for the static case. This provides an introduction to the important metric of makespan, as well as the use of randomization in designing contention-resolution algorithms. Next, in Section 3, we move to dynamic model of packet arrivals, focusing on the case when arrival times are dictated by an adversary. Finally, in Section 4, we address additional adversarial behavior in the form of malicious interference on the channel.
@@
    A discussion portion, where we highlight the main themes of the literature surveyed, along with some potential directions for future work.
@@
    A table that offers a synopsis of the results in each section by organizing the results along common, pertinent attributes from Section 1.1. A brief summary of the result is also provided, which can be useful for reference purposes. Due to space constraints, these tables can be found in our electronic supplement.
@@ 2.1. Backoff Algorithms
In order to better understand why BEB performs poorly, we consider a class of Bernoulli-backoff algorithms, where each packet sends with probability $1/w$ in each slot of a window of size $w$. Note that this differs subtly from the approaches discussed in Section 2.1, where a single slot is chosen uniformly at random in which to send. For Bernoulli-backoff algorithms it is easy to analyze the useful notion of contention (richa:jamming4, ; ChangJP19, ; BenderFGY19, ; DBLP:conf/spaa/AgrawalBFGY20, ; bender:how, ). Let $p_{i}(t)$ denote the probability that packet $i$ attempts to broadcast in slot $t$. We define the contention in slot $t$ to be $C(t)$ $=\sum_{i}p_{i}(t)$; that is, the sum of the sending probabilities for all packets in the system at time $t$.
The notion of contention characterizes the probability that some packet is successful in slot $t$; denote this quantity by $p_{\mbox{\tiny suc}}(t)$. In particular, if $p_{i}(t)\leq 1/2$, which is true if we set the initial window to be size at least $2$, then the following relationship holds:
@@
Like the temperature of the porridge in the Goldilocks fable, the contention can be “too high” or “too low” for packets to succeed. For example, consider the Bernoulli equivalent of BEB. Initially, the contention is high. In particular, $C(t)=\Omega(\ln n)$ which implies that $p_{\mbox{\tiny suc}}(t)$ is polynomially small in $n$ until a window of size $\Theta(n/\ln n)$. We highlight that, while we use Equation 1 to gain intuition about the static case, these bounds on contention will also be useful later on when discussing the dynamic case (Section 3).
@@
Second, from a theoretical perspective, does a backoff algorithm exist that gives optimal throughput under the static case? Bender et al. (bender:adversarial, ) show that, over the class of monotonic windowed backoff algorithms, LLB achieves the asymptotically-best makespan  (bender:adversarial, ). In other words, a non-monotonic approach is required to achieve $\Theta(n)$ makespan or, equivalently, $\Theta(1)$ throughput. This question of asymptotic optimality is what we pursue next.
@@
To develop intuition for monotonic backoff algorithms, we start by asking: why does (the Bernoulli version of) LB fare better than BEB? Again, initially the high contention prevents packets from succeeding. Indeed, the high contention persists for longer under LB given that the window size grows far more slowly compared to BEB. However, once (near) constant contention is achieved, packets succeed. Critically, packets stay within the Goldilocks zone for longer since, again, the window size increases more slowly relative to BEB. Finally, contention drops off more gradually, allowing any straggling packets to succeed within fewer slots than under BEB.
@@
Based on our observations in Section 2, any algorithm that hopes to achieve $O(n)$ makespan must incorporate a mechanism for providing constant contention over $\Omega(n)$ slots. Sawtooth-Backoff (STB) (GreenbergL85, ; Gereb-GrausT92, ) accomplishes this; it is non-monotonic algorithm that executes over a doubly-nested loop. The outer loop sets the current window size $w$ to be double the one used in the preceding outer loop; this is like BEB. Additionally, for each such window, the inner loop executes over $\lg w$ windows of decreasing size: $w,w/2,w/4,...,1$. For each such window, a slot is chosen uniformly at random for the packet to transmit; this is the “backon” component of STB. When the outer loop achieves a window of size $cn$ for a sufficiently large constant $c>0$, with high probability all packets will succeed in the corresponding inner loop. Intuitively, this is because the backon component maintains constant contention in each slot of the inner-loop windows (until all packets succeed). For a single batch of $n$ packets, the makespan of STB is $O(n)$ (Gereb-GrausT92, ; GreenbergL85, ; bender:adversarial, ; zhou2019singletons, ).
A variant called Truncated Sawtooth-Backoff (TSTB) (bender:contention, ) executes identically to STB except that, for some constant $c>0$, the inner loop executes over $\lg(c\lg w)$ windows of decreasing size: $w,w/2,w/4,...,\max\{\lfloor w/c\lg n\rfloor,1\}$. Another asymptotically-optimal result is given by Anta et al. (FernandezAntaMoMu13, ; fernandez-anta:unbounded, ) with their algorithm One-fail Adaptive, which succeeds with probability $1-O(1/n)$, again for the case with no-CD and no knowledge of system size. The authors tightly analyze both One-fail Adaptive and STB, yielding respective makespan guarantees of $2(\delta+1)n+O(\log^{2}n)$ and $4(1+1/\delta)n$, where $\delta$ is a small positive constant.
Anta and Mosteiro (Anta2010, ) look at a formulation of the static case—referred to as static $k$-selection (see Section 3.2)—where an unknown $k$ out of $n$ participants must each succeed. For an error parameter $\epsilon\leq 1/(n+1)$, the authors give a randomized algorithm that with probability $1-O(\epsilon)$ has makespan $O(k+\log^{2}(1/\epsilon))$; for consistency, we report the case for $k=n$ in Table 1 of our electronic supplement.
@@
These results settle the question of optimal makespan for the static case. However, an obvious question is whether this optimality holds when packets may arrive over time, and we explore this issue in Section 3.

### v4

@@ Abstract.
survey, algorithms, contention resolution, backoff, wakeup problem, selection problem, multiple access channel, conflict resolution, Ethernet, adversarial queuing theory, jamming
@@ 1. Introduction
Over the past decade and a half, attention has turned to the case where the arrival times of packets are far-less constrained. One common approach has been to view this problem through the lens of adversarial queuing theory (borodin2001adversarial, ), where results are parameterized by traffic characteristics such as the arrival rate and the degree of burstiness. Another major branch in the literature addresses worst-case arrival times. Arguably, deterministic approaches have been influenced by the pioneering result of Komlós and Greenberg (komlos:asymptotically, ), leading to new solutions for worst-case arrival times. A prime example is the work by De Marco and Kowalski (doi:10.1137/140982763, ), and there has been a long line of results in this area (e.g., (DBLP:conf/icdcs/MarcoKS19, ; Chlebus:2016:SWM:2882263.2882514, ; DEMARCO20171, ; de2013contention, )). In parallel, many different randomized approaches have been explored. A recent influential paper by Chang et al. (ChangJP19, ) shows how the classic ingredients of backing off/on can yield an elegant solution to worst-case arrivals. The literature in these areas is explored in Sections 3.1, 3.2 and 3.3.
@@
Scope and Goals.
In this survey, we focus primarily on contention resolution in the presence of an adversary that (1) controls when participants can begin executing their contention-resolution protocol, and (2) may deny access to the shared resource for periods of time; we refer to this as adversarial contention resolution. This stands in contrast to the setting where arrival times are governed by a stochastic process, which is an area of research that has been previously surveyed by Chlebus (chlebus2001randomized, ).
What motivates our focus on adversarial contention resolution? From a theoretical standpoint, it is interesting to explore what is possible in such a challenging setting, but there are also practical considerations for considering an adversarial context. Network traffic does not always align with stochastic models; specifically, bursty traffic fails to be captured by well-behaved (e.g., Poisson- or Bernoulli-distributed) arrival times. Additionally, the mobility of many modern computing devices, along with connectivity challenges in wireless systems, can lead to situations where the contention for a shared resource changes rapidly over time. Given this motivation, our overall goal in this survey paper is to summarize the models and performance metrics in the literature on adversarial contention resolution, along with discussing their corresponding theoretical results.
@@ 1.1. Definitions, Terminology, Models, Metrics, and Algorithm Design
In the literature, the resource is frequently treated as a shared communication channel, and this is the convention we use throughout this survey. Such a channel is typically referred to as a multiple access channel, which is a popular model under which several fundamental problems have been examined, such as mutual exclusion (bienkowski2010dynamic, ; czyzowicz2009consensus, ), leader election (chang:exponential-jacm, ; chang:exponential, ; nakano2002, ), consensus (czyzowicz2009consensus, ), routing (chlebus2019energy, ), broadcast (Marco08, ; Marco10, ; ChlebusGGPR00, ; alon1991lower, ; ClementiMS03, ; DBLP:conf/podc/Kowalski05, ; chrobak2002fast, ; piotr:explicit, ; chlebus2005almost, ; kowalski2005time, ), and well-known random access protocols (kurose:computer, ). Here, for any given slot, accessing the resource corresponds to a participant listening (equivalently, receiving) or sending (equivalently, transmitting) a packet on the multiple access channel. Unless stated otherwise, a packet can be transmitted within a single slot. When a packet is sent in a slot, we typically say that this packet succeeds if no other packet is sent in the same slot; otherwise, the packet fails.
For ease of exposition, we will often slightly abuse terminology by speaking of packets themselves taking action (e.g., sending themselves on the channel, or listening on the channel). This “packet-centric” presentation allows us to avoid referring to the corresponding “participants”/“devices”/“nodes” that actually perform these actions.
@@
We discuss model assumptions that play an important role in establishing upper and lower bounds with respect to the metrics discussed above.
@@
In contrast, in the dynamic case, packets arrive over time; this arrival model is also referred to in the literature as asynchronous (see  (DBLP:conf/podc/Kowalski05, )). The literature adopts the terminology of arrival time or activation time to indicate when the corresponding participant begins executing a contention-resolution protocol in order to send this packet. These times may be governed by a stochastic process (see the survey by Chlebus (chlebus2001randomized, )), or they may be scheduled by an adversary; in this survey, we focus on the latter setting.
Channel Characteristics. When a participant sends a packet, the multiple access channel provides feedback. The most basic feedback—assumed in all the literature that we survey—occurs when, upon sending its packet, a participant learns whether or not its transmission was successful.
Another type of feedback concerns whether participants can detect collisions. Almost all of the literature on contention resolution assumes one of the following two types of feedback regarding collisions. The first is collision detection (CD): in any slot, each listening participant can differentiate between the case where the number of participants transmitting is 00 (i.e., an empty slot) or greater than $1$ (i.e., a collision). The second is no collision detection (no-CD): in any slot, a listening device cannot differentiate between the slot being empty versus the slot containing a collision.
@@
Finally, some papers consider the case where more than one channel is available to participants; this is typically referred to as a multi-channel setting. Here, typically participants may select any one (and typically only one) channel to send on or listen to in any give slot. A result involving a single channel may extend to the multi-channel setting without altering the core finding; however, there are cases where the use of multiple channels can admit significantly stronger results. Typically, this occurs when multiple channels are leveraged to achieve better latency, or robustness to malicious interference. In practice, multiple channels are available on several communication standards, such as IEEE 802.11 (WiFi), Bluetooth, and cellular standards (e.g., 4G and 5G LTE). Therefore, there is real-world motivation to push for results that make good use of more than one channel. However, in comparison to the single-channel setting, there are fewer results in the contention-resolution literature involving multiple channels.
@@
For the contention-resolution problem, the type of topology may impact channel feedback. With CD in a single-hop network, all participants can detect the same channel feedback in any fixed slot. However, in a multi-hop network, it is common for different sets of participants to witness different channel feedback and, therefore, potentially update their respective states differently. Given this challenge, it is not surprising that the majority of work on contention resolution has focused the single-hop setting.
@@
  • System-Size Knowledge. The exact number of the participants may be known a priori. Alternatively, an upper bound may be common knowledge. In some problem formulations (see Section 3.2), only a subset of $k\leq n$ participants have a packet to send, where $n$ acts as an upper bound that is known or tightly-estimated a priori.
  • Unique Identifiers. Given $n$ participants, each may have a unique identifier (ID) taken from a set of $\Theta(n)$ possible IDs. This ID may be used by an algorithm to determine when the corresponding station accesses the channel.
@@
In the static case, a natural performance measure is the number of slots until a set of packets succeeds; this number of slots is referred to as the makespan. We are typically interested in the set of all packets, but any deviations from this will be made clear from the context. For the static case, the ratio of the number of successful packets to the makespan is known as throughput, and this is another common notion of performance; note that throughput is a value between 00 and $1$.
For the dynamic case, we define a slot to be active, if at least one packet is present in the system during that slot. The above metrics remain valid when measured over active slots. However, at intermediate points in time, throughput may offer an overly-pessimistic measure of progress. For example, if a batch of $n$ packets arrive at the same time and starts executing BEB, the throughput will be 00 for $\Omega(n/\log n)$ slots, despite the fact that the algorithm is progressing towards having many successful packet transmissions, which occurs after $\Theta(n)$ slots are executed.
To provide a more meaningful metric, Bender et al. (bender2020contention, ) propose implicit throughput, which at time $t$ is defined as $N_{t}/A_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$, and $A_{t}$ is the total number of active slots at or before time $t$. Returning to the example with a batch of $n$ packets, we note that the implicit throughput is large—specifically, $n$—in the first slot, and is $\Omega(\log n)$ over the first $\Omega(n/\log n)$ slots. The implicit throughput will continue to decrease, reaching $\Theta(1)$. Then, ultimately, the implicit throughput will be $\Theta(1/\log n)$ once all packets have succeeded, which matches the standard measure of throughput for BEB, as mentioned in Section 1. An algorithm that guarantees an implicit throughout of $\Omega(1)$ over all slots performs well, since the number of active slots so far is never asymptotically larger than the number of packets that have arrived. Furthermore, as exemplified above, at points in time where the standard notion of throughput is meaningful, the two metrics are equal.
Continuing with the dynamic case, another metric is the number of slots it takes for a participant to successfully transmit its packet starting from its activation time. We are interested in the maximum such number of slots over all participants; we refer to this as (maximum) latency . Here, we can extend throughput to be the number of packets over the latency.
@@
Given the popularity of small, battery-powered wireless devices in modern networks, many contemporary results focus on energy efficiency. The energy consumed by a participant is typically measured as the number of slots that the participant accesses the channel, since sending and listening often dominate the operational costs of such devices (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ); we refer to this as the number of channel accesses.
Robustness to malicious interference—often referred to as jamming—is becoming increasingly important, with several well-publicized attacks in recent years (nicholl, ; FCC:humphreys, ; marriott, ; jamming-church, ). Contention-resolution algorithms that rely on channel feedback are especially vulnerable to adversarial noise. We refer to this aspect as jamming resistance and, for each relevant result, we describe the type of attack that can be tolerated.
@@
There is some common terminology used in the literature to characterize different classes of contention-resolution algorithms. An obvious division is with respect to deterministic versus randomized algorithms, where the former is characterized by the absence of random bits being used in the execution. Deterministic solutions are appealing from a practical standpoint in that their guarantees hold even when the number of participants and/or packets is small, admitting no error probability. In trade, they tend to provide sub-optimal asymptotic results in many settings, compared to randomized solutions, as we shall see.
@@ 1.2. Survey Organization
We use the attributes discussed above and illustrated in Figure 1 to structure our review of the literature. In Section 2, we begin with contention-resolution algorithms for the static case. This provides an introduction to the important metric of makespan, as well as the use of randomization in designing contention-resolution algorithms. Next, in Section 3, we move to the dynamic model of packet arrivals, focusing on the case when arrival times are dictated by an adversary; here, latency is a key metric. Finally, in Section 4, we address additional adversarial behavior in the form of jamming.
@@
    A summary portion, where we highlight the main themes of the literature surveyed.
@@
    A table that offers a synopsis of the results in each section by organizing the results along common, pertinent attributes from Section 1.1. Due to space constraints, these tables can be found in our electronic supplement.
@@ 2.1. Backoff Algorithms
In order to better understand why BEB performs poorly, we consider a class of Bernoulli-backoff algorithms (see (bender:adversarial, )), where each packet sends with probability $1/w$ in each slot of a window of size $w$. Note that this differs subtly from the approaches discussed in Section 2.1, where a single slot is chosen uniformly at random in which to send. For Bernoulli-backoff algorithms it is easy to analyze the useful notion of contention (richa:jamming4, ; ChangJP19, ; BenderFGY19, ; DBLP:conf/spaa/AgrawalBFGY20, ; bender:how, ). Let $p_{i}(t)$ denote the probability that packet $i$ attempts to broadcast in slot $t$. We define the contention in slot $t$ to be $C(t)$ $=\sum_{i}p_{i}(t)$; that is, the sum of the sending probabilities for all packets in the system at time $t$.
Contention characterizes the probability that some packet is successful in slot $t$; denote this quantity by $p_{\mbox{\tiny suc}}(t)$. In particular, if $p_{i}(t)\leq 1/2$, which is true if we set the initial window to be size at least $2$, then the following relationship holds:
@@
While we use Equation 1 here to gain intuition about the static case, these bounds on contention will also be useful later on when discussing the dynamic case (Section 3). Like the temperature of the porridge in the Goldilocks fable, the contention can be “too high” or “too low” for packets to succeed. For example, consider the Bernoulli equivalent of BEB. Initially, the contention is high. In particular, $C(t)=\Omega(\ln n)$ until a window of size $\Theta(n/\ln n)$, implying that $p_{\mbox{\tiny suc}}(t)$ is polynomially small (or smaller) in $n$ during this portion of the execution.
@@
Second, from a theoretical perspective, does a backoff algorithm exist that gives optimal throughput under the static case? Bender et al. (bender:adversarial, ) show that, over the class of monotonic windowed backoff algorithms, LLB achieves the asymptotically-best makespan  (bender:adversarial, ). In other words, a non-monotonic approach is required to achieve $\Theta(n)$ makespan or, equivalently, $\Theta(1)$ throughput. This question of asymptotic optimality is what we explore next.
@@
To develop intuition for monotonic backoff algorithms, we start by asking: why does LB fare better than BEB? Considering the Bernoulli versions, initially the high contention prevents packets from succeeding. Indeed, the high contention persists for longer under LB given that the window size grows far more slowly compared to BEB. However, once (near) constant contention is achieved, packets succeed. Critically, packets stay within the Goldilocks zone for longer since, again, the window size increases more slowly relative to BEB. Finally, contention drops off more gradually, allowing any straggling packets to succeed within fewer slots than under BEB.
@@
Based on our observations in Section 2, any algorithm that hopes to achieve $O(n)$ makespan must incorporate a mechanism for providing constant contention over $\Omega(n)$ slots. Sawtooth-Backoff (STB) (GreenbergL85, ; Gereb-GrausT92, ) accomplishes this; it is non-monotonic algorithm that executes over a doubly-nested loop. The outer loop sets the current window size $w$ to be double the one used in the preceding outer loop; this is like BEB. Additionally, for each such window, the inner loop executes over $\lg w$ windows of decreasing size: $w,w/2,w/4,...,1$. For each such window, a slot is chosen uniformly at random for the packet to transmit; this is the “backon” component of STB. When the outer loop achieves a window of size $cn$ for a sufficiently large constant $c>0$, with high probability (i.e., probability at least $1-O(1/{\texttt{poly}({n})})$, an abbreviated as w.h.p.), all packets will succeed in the corresponding inner loop. Intuitively, this is because the backon component maintains constant contention in each slot of the inner-loop windows (until all packets succeed). For a single batch of $n$ packets, the makespan of STB is $O(n)$ (Gereb-GrausT92, ; GreenbergL85, ; bender:adversarial, ; zhou2019singletons, ).
A variant called Truncated Sawtooth-Backoff (TSTB) (bender:contention, ) executes identically to STB except that, for some constant $c>0$, the inner loop executes over $\lg(c\lg w)$ windows of decreasing size: $w,w/2,w/4,...,\max\{\lfloor w/(c\lg w)\rfloor,1\}$. Another asymptotically-optimal result is given by Anta et al. (FernandezAntaMoMu13, ; fernandez-anta:unbounded, ) with their algorithm One-fail Adaptive, which succeeds with probability $1-O(1/n)$, again for the case with no-CD and no knowledge of system size. The authors tightly analyze both One-fail Adaptive and STB, yielding respective makespan guarantees of $2(\delta+1)n+O(\log^{2}n)$ and $4(1+1/\delta)n$, where $\delta$ is a small positive constant.
Anta and Mosteiro (Anta2010, ) look at a formulation of the static case—referred to as static $k$-selection (see Section 3.2)—where an unknown $k$ out of $n$ participants must each succeed. For an error parameter $\epsilon\leq 1/(n+1)$, the authors give a randomized algorithm that with probability $1-O(\epsilon)$ has makespan $O(k+\log^{2}(1/\epsilon))$.
@@
These results settle the question of optimal makespan for the static case using randomized approaches.